# 4D Gaussian Splatting: Modeling Dynamic Scenes with Native 4D Primitives

### v1

@@ VIII-C Ablation and analysis
Our novel approach involves treating 4D Gaussian distributions without strict separation of temporal and spatial elements. In Section III-B, we discussed an intuitive method to extend 3D Gaussians to 4D Gaussians, as expressed in equation (6). This method assumes independence between the spatial $(x,y,z)$ and temporal variable $t$, resulting in a block diagonal covariance matrix. The first three rows and columns of the covariance matrix can be processed similarly to 3D Gaussian splatting. We further additionally incorporate 1D Gaussian to account for the time dimension.
To compare our unconstrained 4D Gaussian with this baseline, we conduct experiments on two representative scenes, as shown in Table III. We can observe the clear superiority of our unconstrained 4D Gaussian over the constrained baseline. This underscores the significance of our unbiased and coherent treatment of both space and time aspects in dynamic scenes. We also qualitatively compare the sliced 3D Gaussians of these two variants in Figure 4. It can be found that under the No-4DRot setting the rim of the wheel is not well reconstructed, and fewer Gaussians are engaged in rendering the displayed frames after filtering, despite a larger total number of fitted Gaussians under this configuration. This indicates that the 4D Gaussian in the No-4DRot setting has less temporal variance, which impairs the capacity of motion fitting and exchanging information between successive frames, and brings more flicker and blur in the rendered videos.
@@
If the 4D Gaussian has only local support in time, as the 3D Gaussian does in space, the number of 4D Gaussians may become very intractable as the video length increases. Fortunately, the anisotropic characteristic of Gaussian offers a prospect of avoiding this predicament. To further unleash the potential of this characteristic, we set the initial time scaling to half of the scene’s duration as mentioned in Section VIII-A2.
@@
More ablations
To examine whether modeling the spatiotemporal evolution of Gaussian’s appearance is helpful, we ablate 4DSH in the second row of Table III. Compared to the result of our default setting, we can find there is indeed a decline in the rendering quality. Moreover, when turning our attention to 4D spacetime, we realize that over-reconstruction may occur in more than just space. Thus, we allow the Gaussian to split in time by sampling new positions using complete 4D Gaussian as PDF. The last two rows in Table III verified the effectiveness of the densification in time.
VIII-D Experiments for the compact 4D Gaussian
VIII-D1 Gaussian attributes compression
TABLE IV: The impact of parameter compression. Metrics are evaluated on the cut roasted beef. All differences are compared to the baseline. For each attribute, we report the compression ratio with respect to the storage of itself.
@@ VIII-D1 Gaussian attributes compression
Implementation details
The baseline model adopts the default settings except for increasing the opacity pruning threshold to 0.05. For scaling, we employ an R-VQ model with $6$ codebooks each containing $64$ entries. The codebook parameters are optimized during $512$ extra steps with a learning rate $4\times 10^{-4}$ on the pre-trained baseline model. Both rotation parameters share the same R-VQ model also consisting of $6$ codebooks with 64 entries, optimized for $1024$ steps at $8\times 10^{-4}$ learning rate. For appearance modeling, the direct and rest components adopt $6$ codebooks with $64$ entries and $8$ codebooks with $256$ entries, respectively. For opacity, we use quantization-aware fine-tuning with $k$-bit Min-Max Quantization [38, 103] to fine-tune for $1000$ iterations with 0.1×0.1$\times$0.1 × learning rate. All experiments are conducted on the Cut Roasted Beef from [4].
Results
In Table IV, we validate the effects of the parameter compression strategy introduced in Section IV-B. For the shape-related attributes, i.e., rotation and scaling, the R-VQ reduces their memory from $12\times 32\times N$ bits to $3\times 6\times 6\times N+2\times 6\times 64\times 4\times 32$ bits for $N$ Gaussians. Since the real scenes could contain millions of Gaussians, this process can achieve around 3.5×3.5$\times$3.5 × compression rates with negligible damage to quality. After VQ and Huffman encoding on the codebook indices, their storage consumption decreases from 88.19 MB to 24.20 MB. Although it leads to a 0.49 dB PSNR drop, the post-finetuning of the codebook entries can recover the PSNR by 0.51 dB. Similarly, the size of color-related parameters decreases by nearly 1000 MB with little PSNR drop. Quantization of position and opacity parameters reduces storage overhead by 50% and 75% respectively, with a negligible performance cost. And the post-finetuning even yields better performance than the original model. By utilizing all of the above strategies together, the total memory overhead decreases from 1183.21 MB to 56.74 MB, while almost maintaining the same rendering quality.
VIII-D2 Culling redundant Gaussians
@@
TABLE V: Impact of pruning thresholds of opacity.
@@
As shown in Figure 8, the total number of Gaussian points decreases rapidly after the stop of densification when using the volume mask, while the PSNR remains stable, indicating that the mask pruning technique effectively removes redundant Gaussians. This results in a 2.47×2.47$\times$2.47 × reduction in the total number of points, from 1.26M to 0.51M. Table V provides the detailed quantitative results. A higher opacity threshold eliminates floaters in the scene, leading to a higher PSNR (33.46 vs. 30.48) and decreasing the number of Gaussian points from 2.62M to 1.26M. Additionally, applying mask pruning further reduces the number of redundant Gaussians while preserving similar rendering quality. Given that the mask pruning technique introduces minimal additional computational overhead, we choose to use it with a threshold of 0.05 as the default method for reducing the number of Gaussians.
VIII-D3 Structured 4D Gaussians
TABLE VI: Ablations on the structured 4D Gaussian.
[…]
Bouncingballs 33.54 0.982 127K
37.78 0.990 5.67K
Lego 25.56 0.938 389K
30.43 0.958 15.6K
In Table VI, we verify the effectiveness of structured 4D Gaussians in reducing the number of primitives and the performance gains in monocular dynamic scenes. Compared to full 4DGS, there is a substantial reduction in the number of final 4D primitives to be stored after optimization, thanks to the reduction of its redundancy in space and time as well as the elimination of a large number of Gaussian caused by overfitting. It also brings considerable performance improvements by better aggregating multi-frame supervision.
VIII-E Experiments on dynamic urban scenes
TABLE VII: Qualitative results of tasks on the dynamic urban scenes. Our results are measured on the Dynamic-32 split of the Waymo-NOTR dataset.
@@ VIII-E Experiments on dynamic urban scenes
VIII-E1 Experimental setup
We evaluate the Waymo-NOTR Dynamic-32 split curated by [61], with 32 sequences of calibrated images captured by five pinhole cameras and LiDAR point clouds. These driving sequences contain rich challenging dynamic objects. Following the common practice, we divide each of them into clips of 50 consecutive frames for training and use images at resolution $640\times 960$ from three frontal cameras in both training and evaluation. For reconstruction, we train and test on all frames of each clip. For novel view synthesis, we take out one frame every ten frames for testing and train on the remaining frames. The warm-up steps for all experiments is set to 3,000 unless otherwise specified.
VIII-E2 Results
Table VII shows that our 4DGS (see Section V) can achieve comparable and even better performance against those complex pipelines specifically designed for urban scenes. Meanwhile, our fully explicit and unified representation supports highly efficient rendering, achieving superior efficiency over all competitors except 3DGS. As shown in Figure 9, 4DGS works well under diverse lighting and weather conditions. It faithfully reconstructs high-frequency texture details and correctly models the geometry for both dynamic and static regions.
VIII-E3 Ablations
TABLE VIII: Ablation studies on the dynamic urban scene. All metrics are evaluated on the first clip of seg104444.
@@ VIII-E3 Ablations
(a) +$\mathcal{L}_{\mathrm{lidar}}$ 36.98 0.960 29.59 0.884
(b) +$\mathcal{L}_{\mathrm{sparse}}$ 36.17 0.961 30.22 0.898
(c) +$\mathcal{L}_{\mathrm{covt}}$ 36.08 0.956 30.49 0.904
@@
(f) +$\mathcal{L}_{\mathrm{rigid}}$ (Full) 36.94 0.959 31.62 0.916
@@
We conduct ablation studies to investigate the impact of modifications for driving scenes. From Table VIII, we can draw these observations: (a) Introducing LiDAR supervision slightly increases performance for novel view synthesis, as relying solely on multi-view images to reconstruct the accurate geometry of textureless regions such as road surfaces is inherently difficult without additional prior constraints. (b) While strong capability of 4DGS, it is challenging to constrain the behavior of Gaussians at a very novel time without static prior. $\mathcal{L}_{sparse}$ effectively addresses this by encouraging sparse motion and significantly boosts the rendering quality in novel views. (c) Applying regularization to the covariance in time dimension can increase the lifetime of the 4D Gaussian, enabling it to better utilize multi-view supervision. (d) Temporal random perturbation further reduces undesired motions, effectively enhancing the performance for novel view synthesis. (e) Warm-up and (f) $\mathcal{L}_{rigid}$ both slightly improve the performance of novel view synthesis by introducing reasonable priors about motion.
VIII-F Experiments for generative 4D Gaussian
TABLE IX: Comparison of 4D representations in video-to-4D generation. To perform fair comparison, we only use anchored images and SDS loss as supervision without any other techniques introduced in representative methods.
[…]
C-DyNeRF [14] Consistent4D [14] 3h (11s) 0.875 0.147
Deform-GS [42] DG4D [73] 4DGen [74] STAG4D [75] 22m (2.6s) 0.898 0.133
4DGS (Ours) Efficient4D [76] 9m30s (1.1s) 0.905 0.129
Figure 10: Segmentation results. We show two views of the scene decomposition and object segments at different scales.
Dataset
Consistent4D test dataset [14]: comprises seven videos with 32 frames, each containing $32\times 4$ ground truth images from four viewpoints at every timestamp. In this dataset, we initialize 50,000 random points inside a sphere of radius 0.5 with uniformly distributed timestamps.
Implementation details
[…]
Competitors
We include cascade dynamic NeRF (C-DyNeRF) introduced in [14] and Deform-GS [42] as alternative representations for comparison. Due to different convergence speeds, we train C-DyNeRF for 1000 iterations and Deform-GS for 500 iterations. The other training settings follow those specified in the respective papers.
Metrics
For each 4D representation, we report the total generation time needed for complete convergence and the time required for a single training iteration. We report the CLIP similarity and LPIPS between rendered images and groundtruth images.
Results
Table IX reports the results on the Consistent4D test dataset. Our 4DGS significantly surpasses all alternatives in terms of both generation time and quality. Figure 11 shows that 4DGS is fastest to converge. We note C-DyNeRF fails to converge in 500 iterations, whereas Gaussian-based representations succeed.
VIII-G Experiments for 4D Gaussian segmentation
We evaluate 4DGS for 4D segmentation (see Section VII) with focus on visualization.
Implementation details
@@ VIII-G Experiments for 4D Gaussian segmentation
We set feature dimensions $D=32$ and $k=16$ in k-nearest neighbors. We train the features and scale mapping based on compact 4DGS (see Section IV), pre-trained on the Sear Steak sequence from the Plenoptic Video dataset. The training is conducted with 10,000 iterations at a batch size of 1. In inference, as [81] we cluster Gaussians in 4D space based on the learned features using HDBSCAN, achieving automatic scene decomposition. For 4D segments, we first extract the 2D feature from the rendered feature map at a user-specific 2D point and then filter the 4D Gaussians whose smoothed scale-gated feature similarity with the query feature is below a certain threshold (typically 0.8). The filtered 4D Gaussians represent a certain component most similar to the 2D point and can be rendered from any views at any timestamp. For both scene decomposition and 4D segments, we adjust the input scale from 0 to 1 to achieve multi-granularity segmentation.
Results
@@
Figure 10 shows the multi-granularity segmentation results. The learned features enable the 4D Gaussians to cluster automatically and adaptively as the scale varies. At scale 0, individual elements like the man’s hair, clothing, and arms are distinctly segmented. As the scale increases to 1, the entire figure of the man is segmented as a whole. This multi-granularity segmentation capability also extends to 4D segments. For example, the Gaussians representing the man’s whole body, head, or arm can be identified at different scales with minimal blurry floaters.
IX Conclusion

### v2

@@ 8.1.3 Ablation and analysis
Our novel approach involves treating 4D Gaussian distributions without strict separation of temporal and spatial elements. In Section 3.2, we discussed an intuitive method to extend 3D Gaussians to 4D Gaussians, as expressed in equation (6). This method assumes independence between the spatial $(x,y,z)$ and temporal variable $t$, resulting in a block diagonal covariance matrix. The first three rows and columns of the covariance matrix can be processed similarly to 3D Gaussian splatting. We further additionally incorporate 1D Gaussian to account for the time dimension.
To compare our unconstrained 4D Gaussian with this baseline, we conduct experiments on two representative scenes, as shown in Table 3. We can observe the clear superiority of our unconstrained 4D Gaussian over the constrained baseline. This underscores the significance of our unbiased and coherent treatment of both space and time aspects in dynamic scenes. We also qualitatively compare the sliced 3D Gaussians of these two variants in Figure 5. It can be found that under the No-4DRot setting the rim of the wheel is not well reconstructed, and fewer Gaussians are engaged in rendering the displayed frames after filtering, despite a larger total number of fitted Gaussians under this configuration. This indicates that the 4D Gaussian in the No-4DRot setting has less temporal variance, which impairs the capacity of motion fitting and exchanging information between successive frames, and brings more flicker and blur in the rendered videos.
@@
If the 4D Gaussian has only local support in time, as the 3D Gaussian does in space, the number of 4D Gaussians may become very intractable as the video length increases. Fortunately, the anisotropic characteristic of Gaussian offers a prospect of avoiding this predicament. To further unleash the potential of this characteristic, we set the initial time scaling to half of the scene’s duration as mentioned in Section 8.1.1.
@@
More ablations To examine whether modeling the spatiotemporal evolution of Gaussian’s appearance is helpful, we ablate 4DSH in the second row of Table 3. Compared to the result of our default setting, we can find there is indeed a decline in the rendering quality. Moreover, when turning our attention to 4D spacetime, we realize that over-reconstruction may occur in more than just space. Thus, we allow the Gaussian to split in time by sampling new positions using a complete 4D Gaussian as PDF. The last two rows in Table 3 verified the effectiveness of the densification in time.
8.2 4D from monocular video
Table 4: Qualitative comparison on monocular dynamic scenes. The LPIPS are computed using VGGNet [110].
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
- D-NeRF (synthetic, monocular)
T-NeRF [9] 29.51 0.95 0.08
D-NeRF [9] 29.67 0.95 0.07
TiNeuVox [108] 32.67 0.97 0.04
HexPlanes [7] 31.04 0.97 0.04
K-Planes-explicit [6] 31.05 0.97 -
K-Planes-hybrid [6] 31.61 0.97 -
V4D [109] 33.72 0.98 0.02
4DGS-HexPlanes [41] [^1] Rendered at $\times800$, otherwise downsampled 2x by default. 33.30 0.98 0.03
4DGS (Ours) 34.09 0.98 0.02
8.2.1 Results of 4D reconstruction
Novel view synthesis for dynamic scenes from monocular video is an inherently under-constrained problem. An ideal solution often requires incorporating proper architectural priors. Although our method, designed as a general-purpose representation for dynamic scenes, does not directly introduce such priors, it still achieves comparable or even better performance than prior art. To validate that, we evaluate our approach on the commonly used monocular dataset. Since the real-world monocular videos for dynamic scenes usually suffer from inaccurate pose estimation, and the effectiveness in complex real-world scenes has been thoroughly validated in the previous section, we test the monocular dynamic reconstruction D-NeRF dataset [9], which comprises 8 object-centric videos synthesized using graphic engine with accurate camera poses accessible. For experiments on this dataset, we initialize the mean position of 4D Gaussians with 100,000 randomly selected points evenly distributed within the cubic volume defined by $1.2,1.2]^{3}$. suggest that the proposed general representation can also work well in such an ill-posed task, highlighting its versatility and potential.
8.2.2 4D from monocular video with generative prior
Table 5: Comparison of 4D representations in video-to-4D generation. To perform fair comparison, we only use anchored images and SDS loss as supervision without any other techniques introduced in representative methods.
[…]
C-DyNeRF [73] Consistent4D [73] 3h (11s) 0.875 0.147
Deform-GS [41] DG4D [76] 4DGen [77] STAG4D [78] 22m (2.6s) 0.898 0.133
4DGS (Ours) Efficient4D [79] 9m30s (1.1s) 0.905 0.129
Figure 10: Segmentation results. We show two views of the scene decomposition and object segments at different scales. Please refer to our project page for the corresponding videos.
Dataset Consistent4D test dataset [73]: comprises seven videos with 32 frames, each containing $32\times 4$ ground truth images from four viewpoints at every timestamp. In this dataset, we initialize 50,000 random points inside a sphere of radius 0.5 with uniformly distributed timestamps.
Implementation details By default, we conduct training with a total of 500 iterations and halted densification at the 100thth iteration. The weights in equation (6) are set as $\lambda_{\text{img}}=10,000$ and $\lambda_{\text{sds}}=1$. The other settings for hyperparameters are the same as reconstruction in Section 8.1.1.
Baselines We include cascade dynamic NeRF (C-DyNeRF) introduced in [73] and Deform-GS [41] as alternative representations for comparison. Due to different convergence speeds, we train C-DyNeRF for 1000 iterations and Deform-GS for 500 iterations. The other training settings follow those specified in the respective papers.
Metrics For each 4D representation, we report the total generation time needed for complete convergence and the time required for a single training iteration. We report the CLIP similarity and LPIPS between rendered images and ground-truth images.
Results Table 5 reports the results on the Consistent4D test dataset. Our 4DGS significantly surpasses all alternatives in terms of both generation time and quality. Figure 9 shows that 4DGS is fastest to converge. We note C-DyNeRF fails to converge in 500 iterations, whereas Gaussian-based representations succeed.
8.3 Experiments for the compact 4D Gaussian
8.3.1 Gaussian attributes compression
Table 6: The impact of parameter compression. Metrics are evaluated on the cut roasted beef. All differences are compared to the baseline. For each attribute, we report the compression ratio with respect to the storage of itself.
@@ 8.3.1 Gaussian attributes compression
Implementation details The baseline model adopts the default settings except for increasing the opacity pruning threshold to 0.05. For scaling, we employ an R-VQ model with $6$ codebooks each containing $64$ entries. The codebook parameters are optimized during $512$ extra steps with a learning rate $4\times 10^{-4}$ on the pre-trained baseline model. Both rotation parameters share the same R-VQ model also consisting of $6$ codebooks with 64 entries, optimized for $1024$ steps at $8\times 10^{-4}$ learning rate. For appearance modeling, the direct and rest components adopt $6$ codebooks with $64$ entries and $8$ codebooks with $256$ entries, respectively. For opacity, we use quantization-aware fine-tuning with $k$-bit Min-Max Quantization [37, 111] to fine-tune for $1000$ iterations with $0.1\times$ learning rate. All experiments are conducted on the Cut Roasted Beef from [4].
Results In Table 6, we validate the effects of the parameter compression strategy introduced in Section 4.2. For the shape-related attributes, i.e., rotation and scaling, the R-VQ reduces their memory from $12\times 32\times N$ bits to $3\times 6\times 6\times N+2\times 6\times 64\times 4\times 32$ bits for $N$ Gaussians. Since the real scenes could contain millions of Gaussians, this process can achieve around $3.5\times$ compression rates with negligible damage to quality. After VQ and Huffman encoding on the codebook indices, their storage consumption decreases from 88.19 MB to 24.20 MB. Although it leads to a 0.49 dB PSNR drop, the post-finetuning of the codebook entries can recover the PSNR by 0.51 dB. Similarly, the size of color-related parameters decreases by nearly 1000 MB with little PSNR drop. Quantization of position and opacity parameters reduces storage overhead by 50% and 75% respectively, with a negligible performance cost. And the post-finetuning even yields better performance than the original model. By utilizing all of the above strategies together, the total memory overhead decreases from 1183.21 MB to 56.74 MB, while almost maintaining the same rendering quality.
8.3.2 Culling redundant Gaussians
@@
Table 7: Impact of pruning thresholds of opacity.
@@
As shown in Figure 11, the total number of Gaussian points decreases rapidly after the stop of densification when using the volume mask, while the PSNR remains stable, indicating that the mask pruning technique effectively removes redundant Gaussians. This results in a $2.47\times$ reduction in the total number of points, from 1.26M to 0.51M. Table 7 provides the detailed quantitative results. A higher opacity threshold eliminates floaters in the scene, leading to a higher PSNR (33.46 vs. 30.48) and decreasing the number of Gaussian points from 2.62M to 1.26M. Additionally, applying mask pruning further reduces the number of redundant Gaussians while preserving similar rendering quality. Given that the mask pruning technique introduces minimal additional computational overhead, we choose to use it with a threshold of 0.05 as the default method for reducing the number of Gaussians.
8.4 Experiments on dynamic urban scenes
Table 8: Qualitative results of tasks on the dynamic urban scenes. Our results are measured on the Dynamic-32 split of the Waymo-NOTR dataset.
@@ 8.4 Experiments on dynamic urban scenes
8.4.1 Experimental setup
We evaluate the Waymo-NOTR Dynamic-32 split curated by [61] from Waymo Open Dataset [114], with 32 sequences of calibrated images captured by five pinhole cameras and LiDAR point clouds. These driving sequences contain rich challenging dynamic objects. We divide each of them into clips of 50 consecutive frames for training and use images at resolution $640\times 960$ from three frontal cameras in both training and evaluation. For reconstruction, we train and test on all frames of each clip. For novel view synthesis, we take out one frame every ten frames for testing and train on the remaining frames. The warm-up steps for all experiments is set to 3,000 unless otherwise specified.
8.4.2 Results
Table 8 shows that our 4DGS (see Section 5) can achieve comparable and even better performance against those complex pipelines specifically designed for urban scenes. Meanwhile, our fully explicit and unified representation supports highly efficient rendering, achieving superior efficiency over all competitors except 3DGS. As shown in Figure 12, 4DGS works well under diverse lighting and weather conditions. It faithfully reconstructs high-frequency texture details and correctly models the geometry for both dynamic and static regions.
8.4.3 Ablations
Table 9: Ablation studies on the dynamic urban scene. All metrics are evaluated on the first clip of seg104444.
@@ 8.4.3 Ablations
(a) $\mathcal{L}_{\mathrm{lidar}}$ 36.98 0.960 29.59 0.884
(b) $\mathcal{L}_{\mathrm{sparse}}$ 36.17 0.961 30.22 0.898
(c) $\mathcal{L}_{\mathrm{covt}}$ 36.08 0.956 30.49 0.904
@@
(f) $\mathcal{L}_{\mathrm{rigid}}$ (Full) 36.94 0.959 31.62 0.916
@@
We conduct ablation studies to investigate the impact of modifications for driving scenes. From Table 9, we can draw these observations: (a) Introducing LiDAR supervision slightly increases performance for novel view synthesis, as relying solely on multi-view images to reconstruct the accurate geometry of textureless regions such as road surfaces is inherently difficult without additional prior constraints. (b) While strong capability of 4DGS, it is challenging to constrain the behavior of Gaussians at a very novel time without static prior. $\mathcal{L}_{sparse}$ effectively addresses this by encouraging sparse motion and significantly boosts the rendering quality in novel views. (c) Applying regularization to the covariance in time dimension can increase the lifetime of the 4D Gaussian, enabling it to better utilize multi-view supervision. (d) Temporal random perturbation further reduces undesired motions, effectively enhancing the performance for novel view synthesis. (e) Warm-up and (f) $\mathcal{L}_{rigid}$ both slightly improve the performance of novel view synthesis by introducing reasonable priors about motion.
8.5 Experiments for 4D Gaussian segmentation
We evaluate 4DGS for 4D segmentation (see Section 7) with focus on visualization.
@@ 8.5 Experiments for 4D Gaussian segmentation
Implementation details We set feature dimensions $D=32$ and $k=16$ in k-nearest neighbors. We train the features and scale mapping based on compact 4DGS (see Section 4), pre-trained on the Sear Steak sequence from the Plenoptic Video dataset. The training is conducted with 10,000 iterations at a batch size of 1. In inference, as [85] we cluster Gaussians in 4D space based on the learned features using HDBSCAN, achieving automatic scene decomposition. For 4D segments, we first extract the 2D feature from the rendered feature map at a user-specific 2D point and then filter the 4D Gaussians whose smoothed scale-gated feature similarity with the query feature is below a certain threshold (typically 0.8). The filtered 4D Gaussians represent a certain component most similar to the 2D point and can be rendered from any views at any timestamp. For both scene decomposition and 4D segments, we adjust the input scale from 0 to 1 to achieve multi-granularity segmentation.
@@
Results Figure 10 shows the multi-granularity segmentation results. The learned features enable the 4D Gaussians to cluster automatically and adaptively as the scale varies. At scale 0, individual elements like the man’s hair, clothing, and arms are distinctly segmented. As the scale increases to 1, the entire figure of the man is segmented as a whole. This multi-granularity segmentation capability also extends to 4D segments. For example, the Gaussians representing the man’s whole body, head, or arm can be identified at different scales with minimal blurry floaters.
9 Conclusion